# Inflationary attractors and radiative corrections in light of ACT data

### v3

@@ III.3 Toy model with matter fields
Fixing $\delta=0.01$, we can then solve for $\kappa$. We then explore for three different choices for all $n$ considered in Fig. 2: (i) for $\xi=1$ and $\lambda\simeq 10^{-10}$ (chosen to match the amplitude of scalar fluctuations observed in the CMB) and (ii) for $\xi=10$ and $\lambda\simeq 10^{-8}$. Furthermore, here $\chi_{\mathrm{cmb}}\simeq 6.0,5.5$ for the choices above. Both of these choices, for all $n$ considered here, result in nearly identical predictions for $(n_{s},r)$ due to their attractor behavior (i.e. for each choice of $n$ at $\xi\simeq\mathcal{O}(1)$ or greater, the predictions are insensitive to increasing $\xi$), but have different implications for the simple toy model with an additional massive field. We can then determine the effective mass $\mathcal{M}^{2}_{\sigma}\simeq\kappa f(\varphi)$ in order to produce $\delta=0.01$ radiative corrections for each member of the $\xi$-attractors.
@@ IV Discussion
Second, there are some signals that should shed light on the viability of the single-field paradigm. For example, single-field models quite generically predict a small negative running of the spectral index $\alpha_{s}\equiv dn_{s}/d\ln k$, where the value lies between $8\times 10^{-3}<\alpha_{\mathrm{S}}<-9.1\times 10^{-5} [11$8]. Radiative corrections do not spoil this for any of the models considered here, so they still are within the scope of this generic prediction. The recent ACT results have also notably shifted the constraints on $\alpha_{s}$ towards the positive direction (see [25, Fig. 4]). While single field inflation models like those considered here are still compatible with the constraints, in addition to the rightward shift in $n_{s}$, this may be another hint that some of the historically more favoured models of inflation may be in tension with new data. If constraints on $\alpha_{s}$ continue to migrate towards the positive direction, this may necessitate moving away from the class of simple single-field inflation altogether. Similarly, investigations relating to generating particular features in the primordial power spectrum, producing primordial black holes, and certain signals in the stochastic gravitational wave background also might also have something to say regarding the ultimate viability of the more basic single-field inflation paradigm when compared with more intricate constructs within single-field modeling, multi-field models, or even more exotic alternatives [119, 120, 121, 122, 123, 124, 125, 126].

### v4

# arXiv:2506.12436v4[astro-ph.CO] 27 Feb 2026

@@ III.3 Toy model with matter fields
Fixing $\delta=0.01$, we can then solve for $\kappa$. We then explore two different choices for all $n$ considered in Fig. 2: (i) for $\xi=1$ and $\lambda\simeq 10^{-10}$ (choosing $\lambda$ to match the amplitude of scalar fluctuations observed in the CMB) and (ii) for $\xi=10$ and $\lambda\simeq 10^{-8}$. Furthermore, here $\chi_{\mathrm{cmb}}\simeq 6.0,5.5$ for the choices above. Both of these choices, for all $n$ considered here, result in nearly identical predictions for $(n_{s},r)$ due to their attractor behavior (i.e. for each choice of $n$ at $\xi\simeq\mathcal{O}(1)$ or greater, the predictions are insensitive to increasing $\xi$), but have different implications for the simple toy model with an additional massive field. We can then determine the effective mass $\mathcal{M}^{2}_{\sigma}\simeq\kappa f(\varphi)$ in order to produce $\delta=0.01$ radiative corrections for each member of the $\xi$-attractors.
@@ IV Discussion
Second, there are some signals that should shed light on the viability of the single-field paradigm. For example, single-field models quite generically predict a small negative running of the spectral index $\alpha_{s}\equiv dn_{s}/d\ln k$, where the value lies between $8\times 10^{-3}<\alpha_{\mathrm{S}}<-9.1\times 10^{-5} [11$8]. Radiative corrections do not spoil this for any of the models considered here, so they still are within the scope of this generic prediction. The recent ACT results have also notably shifted the constraints on $\alpha_{s}$ towards the positive direction (see [25, Fig. 4]). While single field inflation models like those considered here are still compatible with the constraints, in addition to the rightward shift in $n_{s}$, this may be another hint that some of the historically more favoured models of inflation may be in tension with new data. If constraints on $\alpha_{s}$ continue to migrate towards the positive direction, this may necessitate moving away from the class of simple single-field inflation altogether. Similarly, investigations relating to generating particular features in the primordial power spectrum, producing primordial black holes, and certain signals in the stochastic gravitational wave background also might have something to say regarding the ultimate viability of the more basic single-field inflation paradigm when compared with more intricate constructs within single-field modeling, multi-field models, or even more exotic alternatives [119, 120, 121, 122, 123, 124, 125, 126].